by TEI TEI

Project Lead: ThermalTech

03/14/13 – Background checks for A&M and Atkins & Stang are all approved. Two of three for Triton are approved.
03/18/13 – Triton is sending over project schedule for the job today. They are ready to start next week. The pumps ship on Friday. Brian wants to make sure they can drain down the chilled water and condenser water as needed for this work, TTC told them yes but they still have to coordinate with Jackson and Perkins.
03/25/13 – Triton sent TTC a cost sheet for replacing the existing 6″gate valve with a new Butterfly valve. TTC reviewed all work with Greg and Brian from Triton. They made a small piping change in the condenser water connections which TTC approved on site. The triple duty valve will stay in the system and be rebalanced to match the flow of the new CWP-1 being installed. The gate valve can be replaced for under $400 by Triton. Tony needs to give approval to allocate allowance money for this valve. Triton also noted that the pipe sizes on the Public Works documents are incorrect. The second floor piping is actually 4″ not 3″ and the first floor is actually 3″ not 2.5″. The flow and pressure drops are still fine. Triton is to mark up the As-Built drawings and they will be updated at the end of the job.

03/26/13 – Mike Jackson says all demo’d pumps/motors/rails, etc. needs to go. The only motor to stay will be the new motor (spare) that was purchased last year.

04/17/13 – The installed VFD and disconnect is acceptable for use on this project and meets the safety requirements and lock/tagout protocols of Hamilton County.

04/23/13 – A&M will be on site 4/24 to remove the existing DX-9100 expansion modules that presently control the chiller plant. They will then install a new BacNet controller to control the new chiller plant and tie it into the existing BacNet bus with the new VFD’s and the existing BacNet controller that was installed for the boiler project a couple of years ago. This will allow all of the chiller programming to be in one controller adjacent to the chiller. The intention is to have the chiller plant operational by the end of the day 4/24.

Triton sill try to have the balancer on site 4/24 so A&M can set up the controls with the minimum design flow of 110 GPM through the chiller.

04/24/13 – TTC and MJ review controls and mechanical work for pre punchlist work. Chiller is almost ready to be started.

04/26/13 – AHU-4 is low on flow-probably because of trapped air. CWP-1,2 are labeled backwards from dwgs and the head of the system are lower than designed thus increasing the flow readings on the most of the equipment. This is not the final balance report but is shows that we seem to have enough head and flow in the equipment to meet the design conditions. The final report will be included in the final O&M with the final punchlist for this building. Now waiting on controls update from Roger.

04/26/13 – A&M is still working on final checkout, so they did not have the DP sensors up and functional yet. Jim was going to let TTC know what they were reading so the setpoint could be determined. Without the system control, A&M was not able to place it into minimum flow condition (no control over the valves yet and the pump was in hand). Greg was aware as well, but A&M feels there is air in the lines that feed AHU-4 (east unit on the 2nd floor). It’s flow was much lower than the others regardless of where the pump was run. Greg is going to have maintenance look for air vents. Ultimately A&M still needs to verify AHU-4 flow after venting, verify final DP setting, and check flow through the bypass at AHU-1 with all coils closed. This work won’t take long, but needs to be done to ensure everything is ready before A&M schedules.

05/01/13 – The chiller is running and down to one override for now. TTC visited the site and walked the system with Will and talked to Roger on the phone about the programming. The chiller is up and running and with AHU-1,2,3 in automatic. However we need to keep AHU-4 in 85% open mode to meet the minimum flow of the chiller from field testing. From what TTC saw the dP across the chiller needs to be about 1.8 yesterday to keep it running. The mfg. data says a dP across the evaporator barrel only needs to be 0.9 for 116 gpm. TTC left the AHU-4 at 85% open and the dP at 1.8 when he left. This makes the VFD run about 64% speed which is pretty good at least there will be energy to save. On the flow issue there are a couple of items: The first issue is that for some reason there are multiple circuit setters that are not needed and they are installed in a bad location for our needs. The second issue is that we cannot easily replace the Hoffman valves at AHU-4 to bleed the air out of the system. Triton can replace the existing ones above the reception area for $988. TTC would prefer to have the County replace them instead and save the money for the water heater change order. The equipment is only $138 each. Is this a possibility? Triton to send a cost sheet over for review. There is a county meeting to discuss all options.

05/01/13 – The final report for the controls checkout and verification was completed and as follows:

AHU-4 has a 2 1/2″ circuit setter installed vs the 2″ installed at all other units. As the valves were already fully insulated when we arrived, I only cut access to the balancing ports and the position indicator. After trying to bleed air from the system, manipulating the other AHU control valves, and increasing the pump speed, Jim Biele and I were not able to get the coil to meet design. It was at the time I removed insulation to be sure we didn’t have some unseen problem, and found that the valve was in fact a 2 1/2″ valve. Using the proper performance data, we had adequate flow with the system at the original setpoint of 9 PSID.

The setpoint was determined as follows: pump at design flow, all AHU’s commanded to full cooling, all coils verified to be at design flow. The DP at that time was established as the setpoint. Upon completion of the venting and balancing on Monday, both DP sensors were reading the same value with the system at max flow. DP across the chiller at that time was approx 5.4 psid.

The final step was to verify minimum flow in the system with all valves closed. With AHU-1 in bypass (and the small unit in the basement that also has a 3 way control valve), AHU-2,3,4 closed, and the pump in auto 9 psid, we only measured about 55 gpm through the bypass of AHU-1. Jim increased the pump speed to see what would be required to achieve approx 110 gpm at the AHU and we were up to a DP of approx 25 psi before we met minimum flow (and the pump was at 85% at that point). I don’t see how we’re goint to meet the minimum flow requirement without another 3 way valve.

The bypass is piped such that there is a circuit setter (fully open presently) in the bypass line, and then a common return circuit setter downstream of the 3 way control valve. I had to use that common valve to set the flow through the coil. It is throttled to a setting of 14 (out of 100). Fully open, the coil would have too much flow but we would then have greater flow in bypass mode.

05/16/13 – The new bypass valve has been ordered. It’s a two week lead item. A&M should be installing in June and then TTC will commission it after it’s up and running. The plan is to restrict the bypass on AHU-1 at its circuit setter to a minimum and just use the new bypass as it is all we should need.

05/30/13 – CHW Bypass was approved, Triton waiting on change order approval for water heater. Final change order work should start in 2 weeks.

06/19/13 – ThermalTech meet with the Tony and Mike. They have concerns and all change order work is on hold until the bypass valve issue is resolved and working. ThermalTech contacted the contractors and advised them that more work needed to be done prior to installing the water heater.

06/24/13 – ThermalTech visited the site and collected valve positions, space temps, and differential pressure readings for the bypass work. All space temps appear normal.

08/22/13 – TTC met with Tony and Mike to present two options for modified control. Option 1 is a main 3-way valve and Option 2 is to retrofit the system with 3-way valves. We agreed to do Option #2.

08/23/13 – TTC approved 3-way valve installation with Triton. New valves are on order. TTC working with Triton and A&M for schedule and timing.

09/26/13 – Triton and A&M will install valves this week and commission system for operation to finish this project.

10/21/13 – 3-way valves installed. Water heater replacement alternate next. Then this project will be complete.

10/27/13 – Jackson reports all seems to be working very well.

12/13/13 – Startup domestic water heater. Denny & Triton on site for final startup. TTC provided some final installation instructions.

01/24/14 – Final paperwork being received by Triton for closeout. Signed change order came in for records.

01/30/14 – Need certified payroll and final closeout from all three contractors.

02/21/14 – Triton contact for closeout pay applications. TEI approved second to last payapp.

03/25/14 – Triton and TTC settled on May 1 warranty date. Final closeout paperwork enroute.

04/14/14 – Project complete final PW rates being sent to John Nester.